OpenStudy (ntgarrett33):

There are 230.8 million people in the US in 2004, but with a .9% growth each year... How many people will there be in 2005?

sam (.sam.):

(0.9/100)(230.8)=2.1 230.8+2.1=232.9

OpenStudy (ntgarrett33):

Thank You!

OpenStudy (anonymous):

P = 230.8 * 1 .09 = 251.6 mil